OpenStudy (anonymous):

what is equivalent to log (sub 36) 6

Parth (parthkohli):

Okay, temme, 36 to the what power is 6?

OpenStudy (anonymous):

no its log with a little 36 and then 6

Parth (parthkohli):

Yeah.. that's what it means

Parth (parthkohli):

\( \color{Black}{\Rightarrow \log_{36}6 \Longrightarrow \text{36 to the what power is 6? That'd be the answer. } }\)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

my choices are log6, log1/6, log6/log36, or log36/log6

OpenStudy (anonymous):

leah why did you post this on my question!

OpenStudy (anonymous):

post your own this is my question i need mine answered

OpenStudy (anonymous):

i am too!

OpenStudy (anonymous):

o cool

OpenStudy (anonymous):

delete your comment beofre i close the question

Parth (parthkohli):

\( \color{Black}{\Rightarrow \log_xy = \Large {\log_x \over \log_y} }\)

OpenStudy (anonymous):

shut up ok dang dont have to be a aass

OpenStudy (anonymous):

are you kidding? no language go away!

OpenStudy (anonymous):

nd thanks parthkohli i got the answer that way! :)

Parth (parthkohli):

yw :)

OpenStudy (anonymous):

lol thanks honey i happen to be very smart for your info

OpenStudy (anonymous):

can you help me parthkohli??? please???

OpenStudy (anonymous):

no im closing the question! ;) bye

OpenStudy (anonymous):

nooooooo

OpenStudy (anonymous):

im sorry hahahahhha

OpenStudy (anonymous):

closing th questions bye

OpenStudy (anonymous):

parth are you there ?????

Parth (parthkohli):

Yeah I am there. Why?

OpenStudy (anonymous):

she wants you the answer the question she posted as a comment to a tottaly different question!

OpenStudy (anonymous):

\[ \log_{36} 6 = \large \log_{6^2} \normalsize 6 = \frac 1 2 \log_6 6 = \frac 1 2 \]

OpenStudy (anonymous):

so its it 1/6 or what?

OpenStudy (anonymous):

i just got two answers?

OpenStudy (anonymous):

@leah14: Please do not spam in this ongoing discussion. Post your question via Ask the question feature. Also please be respectful to others. -Thanks

OpenStudy (anonymous):

thank you! but foolformath, log36/log6 and log1/6 are both answers

OpenStudy (anonymous):

which is it?

OpenStudy (anonymous):

@FoolForMath sure and respect is my last name sir or mam whatever you are

OpenStudy (anonymous):

The answer is \( \frac 12 \)

OpenStudy (anonymous):

your positive? im nervous i need an A.

OpenStudy (anonymous):

i need an a toooooo

OpenStudy (anonymous):

leah go post your own question! and leave me alone my pictures just fine, so go away!

OpenStudy (anonymous):

the answer was log36/log6

OpenStudy (ledah):

@leah14 I could ask you the same thing

OpenStudy (anonymous):

lol all my life i had to fight ill lay em dead before i let them beat me

OpenStudy (anonymous):

@FoolForMath your a moderator please control leah14 this is harrassment

OpenStudy (ledah):

Then you've already been informed this is NOT your question

OpenStudy (ledah):

Go post your own question instead of harassing others.

OpenStudy (anonymous):

@FoolForMath please moderate this situation please!

OpenStudy (anonymous):

Why are you asking questions during a test?

OpenStudy (ledah):

@leah14 there isn't an excuse for this. So don't try making one :D

OpenStudy (anonymous):

@myininaya your a moderator please help

OpenStudy (anonymous):

guys don't keep tagging mods. You have more than enough looking at it.

OpenStudy (ledah):

@strcaitlyn13 I'd have to agree with becca why are you asking this during a test? o_O

OpenStudy (anonymous):

It has been taken care of.

OpenStudy (anonymous):

its not a test like you can be graded, i just wanted that girl to leave me alone, this i just the example given to show how to work out others. im not on a test. if i was i wouldnt be able to leave the page

OpenStudy (ledah):

Ah. My apologies. It was merely a question. Thank you for clarifying it though

OpenStudy (anonymous):

i understand thank you guys though! :)

OpenStudy (ledah):

Your welcome.